OpenStudy (anonymous):

In an examination if a student looses mark for the wrong answer and gets 1 mark for the correct answer. If he got 30 marks for solving54 questions, then how many questions were correct?

OpenStudy (anonymous):

how many mark lose for wrong answer?

OpenStudy (anonymous):

not mentioned in the question

OpenStudy (anonymous):

I'd say 12 wrong, 42 right.

OpenStudy (anonymous):

@ParthKohli need help

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

yap, 42 right and 12 wrong; if wrong answer deduct 1 mark each

OpenStudy (anonymous):

x - y = 30 x + y = 54 x - 30 = y x + (x - 30) = 54 2x - 30 = 54 2x = 84 x = 42 42 - 30 = 12

Parth (parthkohli):

If there are x right, then 54 - x are wrong. 1(x) + -1(54 - x) = 30 x + (-54 + x) = 30 x - 54 + x = 30 2x - 54 = 30 2x = 84 x = 42 x are right, so 42 are right. 54 - x are wrong, 54 - 42 = 12 are wrong.

Parth (parthkohli):

Get it, @Pratika_Rathore >

OpenStudy (anonymous):

No actually the ans. is not matching it is given as 36 and i don't know that fr mwhere the hell that 36 came...because at first i did the same ...

Parth (parthkohli):

How many marks are deducted on a wrong answer?

OpenStudy (anonymous):

Not mentioned...

Parth (parthkohli):

If the student loses a mark, then he/she loses one mark.

Parth (parthkohli):

(42 * 1) + (12 * -1) 42 - 12 30 Answer is correct..

OpenStudy (anonymous):

if 1 marks is deducted for every three wrong answer, then you can get 36 right answer

Parth (parthkohli):

Or, 1/3 deducted for a wrong one..

OpenStudy (anonymous):

thanx